ly-3x=3 6. Graph and solve: ly=-2 / 3 x 4 A N t - 4 - 2 2 4 . I
Answer #1

Giren equations are 4-3x=3 = h any Solve, 4= 3x+3, 4 =-124-4 - 3x+3 = x (341) * = -4-3 4x=an x=-9 put na-2 in ea y-3(^2) = 3pe 山4-30=3 oa8 洋6 5 4 3 (3,0) 100) Aray] f1234564 89 1002 E A1 -10-9-8-7 (Saya3 a2 of °C-2-3) ? ? /* * * * TF, _ y=-22-4 (316